Assume that random guesses are made for eight multiple choice questions on an SAT​ test, so that there are n=8 ​trials, each with probability of success​ (correct) given by p=0.4. Find the indicated probability for the number of correct answers.




1
Expert's answer
2021-09-30T17:58:46-0400

n = 8

p = 0.4

Let x be the number of correct answers.

P(x = 0) = C(8, 0) 0.40 (0.6)8 = 0.016796

P(x = 1) = C(8, 1) 0.41 (0.6)7 = 0.089580

P(x = 2) = C(8, 2) 0.42 (0.6)6 = 0.209019

P(x = 3) = C(8, 3) 0.43 (0.6)5 = 0.278692

P(x = 4) = C(8, 4) 0.44 (0.6)4 = 0.232243

P(x = 5) = C(8, 5) 0.45 (0.6)3 = 0.123863

P(x = 6) = C(8, 6) 0.46 (0.6)2 = 0.041288

P(x = 7) = C(8, 7) 0.47 (0.6)1 = 0.007864

P(x = 8) = C(8, 8) 0.48 (0.6)0 = 0.000655